# Newton polytopes in cluster algebras and $\tau$-tilting theory

### v1

@@ Abstract.
We prove that the cluster monomials in non-initial cluster variables are uniquely determined by the Newton polytopes of their $F$-polynomials for skew-symmetrizable cluster algebras. Accordingly, we prove that the $\tau$-rigid modules and the left finite multi-simibricks in $\tau$-tilting theory are uniquely determined by their Newton polytopes of these modules.
@@ 1. Introduction
Cluster algebras are a class of commutative algebras with an extra combinatorial structure introduced by Fomin and Zelevinsky [fz_2002]. Such algebras are generated by a special set of generators, called cluster variables, which are grouped into overlapping subsets of fixed size, called clusters. A seed is a pair consisting of a cluster ${\bf x}=(x_{1},\ldots,x_{n})$ and a skew-symmetrizable integer matrix $B$. New seeds can be obtained from a given one by a procedure called mutation. The sets of cluster variables and clusters of a cluster algebra are determined by an initial seed $({\bf x},B)$ and iterative mutations. A cluster monomial is a monomial in cluster variables from the same cluster. Cluster monomials are the central objects to study in cluster algebras.
The $g$-vectors and $F$-polynomials are introduced in [fomin_zelevinsky_2007] to study cluster monomials in terms of the initial seed $({\bf x},B)$. More precisely, each cluster monomial $u$ can be written as
@@
Fomin and Zelevinsky [fomin_zelevinsky_2007] conjectured that different cluster monomials have different $g$-vectors, which has been confirmed by Derksen, Weyman and Zelevinsky [DWZ10] for skew-symmetric cluster algebras and by Gross, Hacking, Keel, and Kontsevich for skew-symmetrizable cluster algebras.
$\tau$-tilting theory was introduced by Adachi, Iyama and Reiten [air_2014], which completes the classic tilting theory from the viewpoint of mutations. Various of the fundamental concepts in cluster algebras, such as clusters/seeds, mutations, $g$-vectors, $F$-polynomials etc., were generalized to $\tau$-tilting theory, cf. [air_2014, fei_2019a]. A dictionary of some concepts between cluster algebras and $\tau$-tilting theory are given in Table 1.
@@
It is known that $\tau$-rigid pairs in $\tau$-tilting theory play the role of cluster monomials in cluster algebras. Adachi, Iyama and Reiten [air_2014] proved that different $\tau$-rigid pairs have different $g$-vectors, which is analogous to the result that different cluster monomials have different $g$-vectors in cluster algebras.
Before we turn to the motivation part of this paper. Let us give the definition of Newton polytopes of polynomials and modules used in this paper.
@@
In order to consider this question, it is natural to exclude the initial cluster variables $x_{1},\ldots,x_{n}$ and correspondingly the initial indecomposable $\tau$-rigid pairs $(0,P_{1}),\ldots,(0,P_{n})$, because the $F$-polynomials of these initial objects are $1$.
In [ckq_2024], Keller, Qin and the author of the present paper proved that cluster monomials in non-initial cluster variables are uniquely determined by their $F$-polynomials. The proof there depends on a notion of valuation pairing on (upper) cluster algebras. Typically, the same proof does not work in $\tau$-tilting theory.
Recently, it is proved [cao-2025b] that the indecomposable $\tau$-rigid modules and left finite bricks in $\operatorname{\mathsf{mod}}\nolimits A$ are uniquely determined by their Newton polytopes, which is equivalent to saying that such modules are uniquely determined by the Newton polytopes of their $F$-polynomials, thanks to Fei’s result [fei_2019a]. The proof uses two key results: (i) all functorially finite torsion classes are semistable torsion classes [asai-2021]; and (ii) the brick-$\tau$-rigid correspondence in [DIJ-17], which is a bijection between the indecomposable $\tau$-rigid modules and the left finite bricks. In the general case, it is proved that if two rigid modules $U$ and $V$ (not necessarily indecomposable) have the same Newton polytope, then $U\oplus V$ remains $\tau$-rigid. Unfortunately, it is not clear from [cao-2025b] whether $U$ and $V$ are isomorphic or not. The main motivation for this paper is the following question.
@@
Whether the Newton polytopes of $F$-polynomials determine the cluster monomials in non-initial cluster variables and $\tau$-rigid modules?
Recently, the Newton polytopes of $F$-polynomials and modules were extensively studied in the literature, cf. [BK-2012, BKT-2014, BCDMTY-2024, fei_2019a, fei_2019b, AHIKM-2022, Asai-Iyama-2020, li-pan-2022, cao-2025b]. In particular, the Newton polytopes of modules over preprojective algebras are used to study MV polytopes in [BKT-2014], which parametrize Lusztig’s canonical basis, cf. [Kamnitzer-2010].
@@
(iii) A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is left finite, if the smallest torsion class containing $M$ is a functorially finite torsion class in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
The left finite multi-semibricks are the dual side of $\tau$-rigid modules in some sense. So it is natural to extend the result on $\tau$-rigid modules to the left finite multi-semibricks.
@@
The proofs of Theorem 1.4 and Theorem 1.5 are quite similar, whereas the proof of Theorem 1.6 is slightly different. Let make some comments about the proofs of Theorem 1.4 and Theorem 1.5. Both proofs are based on the reduction arguments. The key points are summarized as follows:
@@
    The proof of Theorem 1.5 is by reducing the number $|U\oplus V|$ of iso-classes of indecomposable direct summands of $U\oplus V$. The proof of Theorem 1.4 is by reducing the number $|\operatorname{\mathsf{supp}}(uv)|$ of the support set $\operatorname{\mathsf{supp}}(uv)$ of $uv$, which is the set of cluster variables appearing in the cluster monomial uvuv.
@@
Of course, many preparations are required to ensure that each stage goes smoothly. These will be detailed in the main body of the paper. Among these, the $F$-invariant and partial $F$-invariant in cluster algebra and $\tau$-tilting theory will play a crucial role in the proofs, because such invariants are directly related to the Newton polytopes. We give a detailed discussion on $F$-invariant of cluster monomials in Subsection 4.1 and that of decorated modules in Subsection 5.1.
This paper contains two closely related parts, cluster algebras and $\tau$-tilting theory. These two parts can be read in any order. Personally, I first proved the main result in cluster algebras and then realized that a similar approach can be applied to the $\tau$-tilting theory.
@@
The prefixes “left” and “right” in the terminology “left mutation”, “right mutation”, “left Bongartz completion”, and “right Bongartz completion” indicate the direction of change under a suitable partial order: “left” denotes a move to a smaller element, while “right” denotes a move to a larger one. This convention aligns with the familiar ordering of integers $(\dots,-2,-1,0,1,2,\ldots)$, where moving left yields smaller numbers and moving right yields larger ones.
@@ 2.1. Polytopes, Minkowski sum and tropical polynomials
In this paper, we mainly focus on the polytopes defined from polynomials and modules of a finite dimension algebra $A$.
@@
Tropical polynomials play an important role in defining the $F$-invariant in cluster algebras and $\tau$-tilting theory in Subsections 4.1, 5.1. We can see that if ${\bf r}\in\mathbb{Z}^{n}$, then $F[{\bf r}]\in\mathbb{Z}$.
@@
Note that if the polynomial $F$ has constant term $1$, then $F[{\bf r}]\in\mathbb{Z}_{\geq 0}$ for any ${\bf r}\in\mathbb{Z}^{n}$. We remark that the $F$-polynomials of cluster monomials and modules always have constant term $1$.
@@ 2.2. Cluster algebras
Let $\check{\mathcal{A}}$ be the cluster algebra with initial seed $({\bf z}_{t_{0}},B_{t_{0}}^{T})$ and we write $({\bf x}_{t},B_{t}^{T})$ for its seed at vertex $t$. We use $\check{G}_{t_{2}}^{t_{1}}$ to denote the $G$-matrix of $({\bf z}_{t_{2}},B_{t_{2}}^{T})$ with respect to vertex $t_{1}$. By [NZ12, (1.13)], we have $C_{t}^{t_{0}}=(\check{G}_{t_{0}}^{t})^{T}$. Since $C_{t}^{t_{0}}$ is a non-negative matrix, we know that $\check{G}_{t_{0}}^{t}$ is a non-negative matrix. Now we take $({\bf z}_{t},B_{t}^{T})$ as the initial seed of $\check{\mathcal{A}}$. The non-negative $G$-matrix $\check{G}_{t_{0}}^{t}$ corresponds to the positive chamber in the scattering diagram [GHKK18] of $\check{\mathcal{A}}$. This implies ${\bf z}_{t_{0}}]=[{\bf z}_{t}]$. So $\check{G}_{t_{0}}^{t}$ is a permutation matrix. Thus $C_{t}^{t_{0}}$ is also a permutation matrix. Therefore, $(C_{t}^{t_{0}})^{-1}$ is a permutation matrix. In particular, it is a non-negative matrix. Then by Theorem 2.18 (i), we see that the $G$-matrix $G_{t}^{t_{0}}=D(C_{t}^{t_{0}})^{-1}D^{-1}$ of $\mathcal{A}$ is a non-negative matrix. So it corresponds to the positive chamber in the scattering diagram of $\mathcal{A}$. Thus ${\bf x}_{t}]=[{\bf x}_{t_{0}}]$. ∎
@@ 2.3. Bongartz completion in cluster algebras
The dominant sets are introduced in [Cao-2023] as a replacement of torsion classes in the $\tau$-tilting theory. The non-initial cluster variables in the dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ correspond to the indecomposable $\tau$-rigid modules contained in the torsion class $\operatorname{\mathsf{Fac}}M=\prescript{\bot}{}{(}{\tau M})\cap P^{\bot}$ for a $\tau$-tilting pair $(M,P)$ in the $\tau$-tilting theory, cf. [Cao-2023, Proposition 7.23] or Corollary 5.11 in this paper.
@@ 3.1. Decorated modules and $\tau$-tilting pairs
We fix a finite dimensional basic algebra $A$ over an algebraically closed field $\mathbf{k}$. Denote by $\operatorname{\mathsf{mod}}\nolimits A$ the category of finitely generated left $A$-modules, and by $\tau$ the Auslander-Reiten translation in $\operatorname{\mathsf{mod}}\nolimits A$. The isomorphism classes of indecomposable projective modules and simple modules in $\operatorname{\mathsf{mod}}\nolimits A$ are denoted by $P_{1},\ldots,P_{n}$.
@@
  • $\operatorname{\mathsf{Fac}}M$ the factor modules of the modules in $\operatorname{\mathsf{add}}\nolimits M$.
  • $\operatorname{\mathsf{Sub}}M$ the submodules of the modules in $\operatorname{\mathsf{add}}\nolimits M$.
@@
Note that the negative parts of decorated modules in our definition are projective $A$-modules, whereas the negative parts of decorated modules using in [DWZ10] are semisimple $A$-modules. Clearly, these two types of decorated $A$-modules are in bijection with each other.
@@
A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is $\tau$-rigid, if $\operatorname{\mathsf{Hom}}_{A}(M,\tau M)=0$.
@@ 3.2. $g$-vectors and $F$-polynomials of decorated modules
Since the zero module is a quotient module of $M\in\operatorname{\mathsf{mod}}\nolimits A$, the polynomial $F_{M}$ has constant term $1$. Similarly, the dual $F$-polynomial $\check{F}_{M}$ also has the constant term $1$. Thus we have
@@
By Propositon 3.8 and Propositon 2.4, we know that
@@
in $\operatorname{\mathsf{mod}}\nolimits A=\operatorname{\mathsf{add}}\nolimits(P_{1}\oplus P_{2}\oplus S_{1})$. The Hasse quiver of the poset of functorially torsion classes in $\operatorname{\mathsf{mod}}\nolimits A$ is as follows:
${\mathcal{T}_{0}=\{0\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{1}=\operatorname{\mathsf{add}}\nolimits(S_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{4}=\operatorname{\mathsf{add}}\nolimits(P_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{2}=\operatorname{\mathsf{add}}\nolimits(S_{1}\oplus P_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{3}=\operatorname{\mathsf{mod}}\nolimits A}$
The corresponding basic $\tau$-tilting pairs are given as follows:

### v2

@@ Abstract.
We prove that the cluster monomials in non-initial cluster variables are uniquely determined by the Newton polytopes of their $F$-polynomials for skew-symmetrizable cluster algebras. Accordingly, we prove that the $\tau$-rigid modules and the left finite multi-semibricks in $\tau$-tilting theory are uniquely determined by the Newton polytopes of these modules. The key tools used in the proofs are the left Bongartz completion, $F$-invariant and partial $F$-invariant in the context of cluster algebras and $\tau$-tilting theory.
@@ 1. Introduction
Cluster algebras are a class of commutative algebras equipped with an extra combinatorial structure introduced by Fomin and Zelevinsky [fz_2002]. Such algebras are generated by a special set of generators, called cluster variables, which are grouped into overlapping subsets of fixed size, called clusters. A seed is a pair consisting of a cluster ${\bf x}=(x_{1},\ldots,x_{n})$ and a skew-symmetrizable integer matrix $B$. New seeds can be obtained from a given one by a procedure called mutation. The sets of cluster variables and clusters of a cluster algebra are determined by an initial seed $({\bf x},B)$ and iterative mutations. A cluster monomial is a monomial in cluster variables from the same cluster. Cluster monomials are the central objects to study in cluster algebras.
The $g$-vectors and $F$-polynomials are introduced in [fomin_zelevinsky_2007] to study the cluster monomials of a cluster algebra in terms of the initial seed $({\bf x},B)$. More precisely, each cluster monomial $u$ can be written as
@@
Fomin and Zelevinsky [fomin_zelevinsky_2007] conjectured that different cluster monomials have different $g$-vectors, which has been confirmed by Derksen, Weyman, and Zelevinsky [DWZ10] for skew-symmetric cluster algebras and by Gross, Hacking, Keel, and Kontsevich for skew-symmetrizable cluster algebras.
$\tau$-tilting theory was introduced by Adachi, Iyama, and Reiten [air_2014], which completes the classic tilting theory from the viewpoint of mutations. Various of the fundamental concepts in cluster algebras, such as clusters/seeds, mutations, $g$-vectors, $F$-polynomials etc., were generalized to $\tau$-tilting theory, cf. [air_2014, fei_2019a, CGY-2023, Cao-2023]. Table 1 summarizes the correspondence between cluster algebras and $\tau$-tilting theory.
@@
It is known that $\tau$-rigid pairs in $\tau$-tilting theory play the role of cluster monomials in cluster algebras. Adachi, Iyama, and Reiten [air_2014] proved that different $\tau$-rigid pairs have different $g$-vectors, which is analogous to the result that different cluster monomials have different $g$-vectors in cluster algebras.
Before discussing the motivation of this paper, we recall the definition of Newton polytopes for polynomials and modules, which will be used throughout. Throughout, let $A$ be a finite dimensional basic algebra over an algebraically closed field $\mathbf{k}$.
@@
In order to consider this question, it is natural to exclude the initial cluster variables $x_{1},\ldots,x_{n}$ and correspondingly the initial indecomposable $\tau$-rigid pairs $(0,P_{1}),\ldots,(0,P_{n})$, because the $F$-polynomials of these initial objects are $1$. So we only need to consider cluster monomials in non-initial cluster variables and $\tau$-rigid modules.
In [ckq_2024], Keller, Qin, and the author proved that the cluster monomials in non-initial cluster variables are uniquely determined by their $F$-polynomials. The proof there depends on a notion of valuation pairing on (upper) cluster algebras. Typically, the same proof does not work in $\tau$-tilting theory.
Recently, it was proved in [cao-2025b] that the indecomposable $\tau$-rigid modules and the left finite bricks in $\operatorname{\mathsf{mod}}\nolimits A$ are uniquely determined by their Newton polytopes, which is equivalent to saying that such modules are uniquely determined by the Newton polytopes of their $F$-polynomials, thanks to Fei’s result [fei_2019a]. The proof in [cao-2025b] is based on two key results: (i) all functorially finite torsion classes are semistable torsion classes [asai-2021]; and (ii) the brick-$\tau$-rigid correspondence in [DIJ-17], which is a bijection between the indecomposable $\tau$-rigid modules and the left finite bricks. In the general case, it is proved in [cao-2025b] that if two rigid modules $U$ and $V$ (not necessarily indecomposable) have the same Newton polytope, then $U\oplus V$ remains $\tau$-rigid (see [cao-2025b, Theorem 3.2] or Lemma 5.12 in this paper). However, it is not clear whether $U$ and $V$ are isomorphic or not. The main motivation for this paper is the following question.
@@
Whether the Newton polytopes of $F$-polynomials determine the cluster monomials in non-initial cluster variables and the $\tau$-rigid modules?
Recently, the Newton polytopes of $F$-polynomials and modules were extensively studied in the literature, cf. [BK-2012, BKT-2014, BCDMTY-2024, fei_2019a, fei_2019b, AHIKM-2022, Asai-Iyama-2020, li-pan-2022, cao-2025b]. In particular, the Newton polytopes of modules over preprojective algebras are used to study MV polytopes in [BK-2012, BKT-2014], which parametrize Lusztig’s canonical basis, cf. [Kamnitzer-2010].
@@
(iii) A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is called left finite, if the smallest torsion class containing $M$ is a functorially finite torsion class in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
The left finite multi-semibricks are the dual counterpart of $\tau$-rigid modules in some sense. So it is natural to extend the result on $\tau$-rigid modules to the left finite multi-semibricks.
@@
The proofs of Theorem 1.4 and Theorem 1.5 are quite similar, while the proof of Theorem 1.6 is slightly different. Let us make some comments about the proofs of Theorem 1.4 and Theorem 1.5. Both proofs are based on the reduction arguments. The key points are summarized as follows:
@@
    The proof of Theorem 1.5 is by reducing the number $|U\oplus V|$ of iso-classes of indecomposable direct summands of $U\oplus V$. The proof of Theorem 1.4 is by reducing the number $|\operatorname{\mathsf{supp}}(uv)|$ of cluster variables in the support set $\operatorname{\mathsf{supp}}(uv)$ of $uv$, which is the set of cluster variables appearing in the cluster monomial uvuv.
@@
Of course, many preparations are required to ensure that each stage goes smoothly. These will be detailed in the main body of the paper. Among these, the $F$-invariant and partial $F$-invariant in cluster algebra and $\tau$-tilting theory will play a crucial role in the proofs, because such invariants are directly related to the Newton polytopes. We give a detailed discussion on $F$-invariant of cluster monomials in Section 4.1 and that of decorated modules in Section 5.1.
This paper consists of two closely related parts: cluster algebras and $\tau$-tilting theory. These two parts can be read in any order. Personally, I first proved the main result in cluster algebras and then realized that a similar approach can be applied to the $\tau$-tilting theory.
@@
In this paper, the prefixes “left” and “right” in the terminology “left mutation”, “left Bongartz completion”, “right mutation”, and “right Bongartz completion” indicate the direction of change under a suitable partial order: “left” denotes a move to a smaller element, while “right” denotes a move to a larger one. This convention aligns with the familiar ordering of integers $(\dots,-2,-1,0,1,2,\ldots)$, where moving left yields smaller numbers and moving right yields larger ones.
@@ 2.1. Polytopes, Minkowski sum and tropical polynomials
In this paper, we mainly focus on the polytopes defined from polynomials and modules of a finite dimensional algebra $A$.
@@
Tropical polynomials play an important role in defining the $F$-invariant in cluster algebras and $\tau$-tilting theory in Sections 4.1, 5.1. We can see that if ${\bf r}\in\mathbb{Z}^{n}$, then $F[{\bf r}]\in\mathbb{Z}$.
@@
We have the following important facts.
  • (a)
    The tropical polynomial $F[-]:\mathbb{R}^{n}\to\mathbb{R}$ is uniquely determined by the Newton polytope $\mathsf{P}(F)$ of $F$. Actually, it only depends on the vertices of the Newton polytope $\mathsf{P}(F)$.
  • (b)
    If ${\bf r}\in\mathbb{Z}^{n}$, then $F[{\bf r}]\in\mathbb{Z}$.
  • (c)
    If the polynomial $F$ has constant term $1$, then $F[{\bf r}]\in\mathbb{Z}_{\geq 0}$ for any ${\bf r}\in\mathbb{Z}^{n}$.
  • (d)
    The $F$-polynomials of cluster monomials and modules defined later always have constant term 11 (see Theorem 2.18 (iv) and Remark 3.8).
@@ 2.2. Cluster algebras
Let $\check{\mathcal{A}}$ be the cluster algebra with initial seed $({\bf z}_{t_{0}},B_{t_{0}}^{T})$ and we write $({\bf z}_{t},B_{t}^{T})$ for its seed at vertex $t$. We use $\check{G}_{t_{2}}^{t_{1}}$ to denote the $G$-matrix of $({\bf z}_{t_{2}},B_{t_{2}}^{T})$ with respect to vertex $t_{1}$. By [NZ12, (1.13)], we have $C_{t}^{t_{0}}=(\check{G}_{t_{0}}^{t})^{T}$. Since $C_{t}^{t_{0}}$ is a non-negative matrix, we know that $\check{G}_{t_{0}}^{t}$ is a non-negative matrix. Now we take $({\bf z}_{t},B_{t}^{T})$ as the initial seed of $\check{\mathcal{A}}$. The non-negative $G$-matrix $\check{G}_{t_{0}}^{t}$ corresponds to the positive chamber in the scattering diagram [GHKK18] of $\check{\mathcal{A}}$. This implies ${\bf z}_{t_{0}}]=[{\bf z}_{t}]$. So $\check{G}_{t_{0}}^{t}$ is a permutation matrix. Thus $C_{t}^{t_{0}}$ is also a permutation matrix. Therefore, $(C_{t}^{t_{0}})^{-1}$ is a permutation matrix. In particular, it is a non-negative matrix. Then by Theorem 2.18 (i), we see that the $G$-matrix $G_{t}^{t_{0}}=D(C_{t}^{t_{0}})^{-1}D^{-1}$ of $\mathcal{A}$ is a non-negative matrix. So it corresponds to the positive chamber in the scattering diagram of $\mathcal{A}$. Thus ${\bf x}_{t}]=[{\bf x}_{t_{0}}]$. ∎
@@ 2.3. Bongartz completion in cluster algebras
Although the remainder of this subsection is not used in the rest of the paper, it is included to provide an analogue of Proposition 3.14 in $\tau$-tilting theory.
@@
The dominant sets are introduced in [Cao-2023] as a replacement of the torsion classes in $\tau$-tilting theory. The non-initial cluster variables in the dominant set $\operatorname{\mathsf{dom}}^{t_{0}}[t]$ correspond to the indecomposable $\tau$-rigid modules contained in the torsion class $\operatorname{\mathsf{Fac}}M=\prescript{\bot}{}{(}{\tau M})\cap P^{\bot}$ for a $\tau$-tilting pair $(M,P)$ in $\tau$-tilting theory, cf. [Cao-2023, Proposition 7.23] or Corollary 5.11 in this paper.
@@ 3.1. Decorated modules and $\tau$-tilting pairs
We fix a finite dimensional basic algebra $A$ over an algebraically closed field $\mathbf{k}$. Denote by $\operatorname{\mathsf{mod}}\nolimits A$ the category of finitely generated left $A$-modules, and by $\tau$ the Auslander-Reiten translation in $\operatorname{\mathsf{mod}}\nolimits A$. The isomorphism classes of indecomposable projective modules in $\operatorname{\mathsf{mod}}\nolimits A$ are denoted by $P_{1},\ldots,P_{n}$.
@@
  • $\operatorname{\mathsf{Fac}}M$ the subcategory of $\operatorname{\mathsf{mod}}\nolimits A$ consisting of the quotient modules of the modules in $\operatorname{\mathsf{add}}\nolimits M$.
  • $\operatorname{\mathsf{Sub}}M$ the subcategory of $\operatorname{\mathsf{mod}}\nolimits A$ consisting of the submodules of the modules in $\operatorname{\mathsf{add}}\nolimits M$.
@@
Note that the negative parts of decorated modules in our definition are projective $A$-modules, whereas the negative parts of decorated modules used in [DWZ10] are semisimple $A$-modules. Clearly, these two types of decorated $A$-modules are in bijection with each other.
@@
A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is called $\tau$-rigid, if $\operatorname{\mathsf{Hom}}_{A}(M,\tau M)=0$.
@@ 3.2. $g$-vectors and $F$-polynomials of decorated modules
Remark 3.8.
Since the zero module is a quotient module of $M\in\operatorname{\mathsf{mod}}\nolimits A$, the polynomial $F_{M}$ has constant term $1$. Similarly, the dual $F$-polynomial $\check{F}_{M}$ also has constant term $1$. Thus we have
@@
By Propositon 3.9 and Propositon 2.4, we know that
@@
in $\operatorname{\mathsf{mod}}\nolimits A=\operatorname{\mathsf{add}}\nolimits(P_{1}\oplus P_{2}\oplus S_{1})$. There are $5$ basic $\tau$-tilting pairs in $\operatorname{\mathsf{mod}}\nolimits A$ given as follows: